Discrete Math

1. Use the primes P1 = 3 and P2 = 17 and the value E = 3 with the RSA algorithm to compute the values below needed to get the keys:

  1. What is the value of N?
  2. What is the value of Z?
  3. What is the value of D?

(ABET 5) Using the values from the problem 9) above, show how to encrypt the message, M = 44. Do NOT simplify, just show the computation needed to encrypt it.

2. (ABET 5) Using the values from the problem 9) above, show how you would decrypt the cipher message, C = 25. Do NOT simplify, just show the computation needed to decrypt it.

3. (ABET 5) Suppose that Jerry Jones, the owner of the Dallas Cowboys, wants to send a message, M, to all the players. But he wants to digitally sign his message so that the players know it was from him.

4. Assume that Jerry Jones has the public key N and E, and the private key, D. Show what Jerry would compute, C, to digitally sign his message, M.
Answer #1

Solutions page . . Given that use the poines P1=3 and and o the value E=3 with RSA algorithm In ouder to solve any RSA algoouPage prime numbers are Given that two and E= 3 P123, P2 = 17 The value of N Ni Pi* P2 1. = 3 * 17 N: 51 value of z z = (1,-1)* Represent the alarn text message as Seo пі% 12мг N possitive integen r with * compute the cipher text = S c HE modn Therefo. let Public - page ② us assume that jenay jones has the key Ni and E. and the private key резонот - Criven n a a cipheatent